OpenStudy (anonymous):

f(x)= 1/x simplify: f(a+h) - f(a) ___________ h I got h/a(a+h) but I'm not sure if I am right. please help if you can :)

OpenStudy (anonymous):

f(a+h) = 1/(a+h) f(a) = 1/a so the difference quotient in your problem is: 1/h * (1/(a+h) - 1/a) get common denominators there..i don't think you'll end up with an h at the top as per your answer.

OpenStudy (lgbasallote):

\[\large \frac{\frac{1}{a+ h} + \frac{1}{a}}{h}\] lol im just latexing

OpenStudy (apoorvk):

f(a+h) = 1/(a+h) f(h) = 1/a so, what you ask for is: \[\large \frac {\frac1{a+h} -\frac{1}{a}} {h}\]

OpenStudy (anonymous):

Yes that @apoorvk

OpenStudy (anonymous):

I did that and found the common denominators but I'm not sure if I got the right answer.

OpenStudy (anonymous):

but my professor wants us to simplify that even more by finding the common denominators and going even further into it.. so tedious -.-

OpenStudy (anonymous):

latex people....^^^

OpenStudy (lgbasallote):

lol it seems i accidentally posted an answer :S against my principles psh psh

OpenStudy (apoorvk):

Did you do this? actually you should get: \[\frac 1 {a(a+h)}\] @dpaInc check your solution once - i seem to find an error.

OpenStudy (anonymous):

well what was the common denominator? I used ah(a+h)

OpenStudy (anonymous):

i didn't post a solution...

OpenStudy (anonymous):

just what sheep needed to simplify...

OpenStudy (anonymous):

she...

OpenStudy (apoorvk):

how are you getting an 'ah(a+h)' in the denominator? you should get only a(a+h) - infact you should also get a 'negative' sign which I forgot to include.

OpenStudy (apoorvk):

@dpaInc I mean there seems to be a sign error - seems like you copied down the question incorrectly

OpenStudy (anonymous):

i'm pretty sure what i wrote as the the difference quotient is correct... it's iggy that had a sign error... (sorry iggy.)

OpenStudy (apoorvk):

\[\large \frac {\frac1{a+h} -\frac{1}{a}} {h}\] \[\large \frac {\frac{a-(a +h)}{a(a+h)}} h\]

OpenStudy (callisto):

I hate chrome!!! When I almost finished typing, snaps happened... \[\frac{f(a+h) - f(a)}{h} = \frac{\frac{1}{(a+h)} - \frac{1}{(a)}}{h} = \frac{\frac{a}{a(a+h)} - \frac{(a+h)}{(a(a+h))}}{h}\]\[= \frac{\frac{a-a-h}{a(a+h)} }{h} = \frac{-h}{ah(a+h)} = -\frac{1}{a(a+h)} \] It looks like differentiate an expression using the first principle :|

OpenStudy (anonymous):

everything is being trumped now.... ^^^

OpenStudy (callisto):

huh?!

OpenStudy (anonymous):

Thank you all for the help!

OpenStudy (apoorvk):

@Callisto switch to Firefox for OS.